Two parallel long wires carry i1 and i2 with i1 greater than i2. when current are in the same direction, the magnetic field midways between the wires is 10 F. when direction of current is reversed it becomes 40F.find ratio of i1/i2.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Let the mid distance be at r.
Then net Magnetic field due to both the wires at that midpoint would be:
Bnet = 10 = k(i1 – i2) 
After the directions of the current is reversed then the net Magnetic field becomes:
Bnet = 40 = k(i2 – i1)
Solving both the equations by dividing we get,
(i1 – i2)/(i2 – i1) ¼
i1/i2  = 1/1 => 1:1